OpenStudy (anonymous):

relative rate of change for f(x)= 5x - xlnx

OpenStudy (anonymous):

Same deal as before. From the definition,\[\frac{f'(x)}{f(x)}=\frac{5-(\ln x+1)}{5x-x \ln x}=\frac{4-\ln x}{x(5-\ln x)}\]

OpenStudy (anonymous):

Are you having trouble with taking the derivative?

OpenStudy (anonymous):

yeah i wasnt sure how to get the derivative.

OpenStudy (anonymous):

jsut when you have ln

OpenStudy (anonymous):

I'm thinking it's the xlnx bit?

OpenStudy (anonymous):

\[\frac{d}{dx}x \ln x=x \frac{d}{dx}\ln x+\ln x \frac{d}{dx}x=x \times \frac{1}{x}+\ln x \times 1\]\[=1+\ln x\]

OpenStudy (anonymous):

So the entire derivative in the numerator is\[f'(x)=5-(1+\ln x)=4-\ln x\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

im trying to correct this test and this class is killing me

OpenStudy (anonymous):

we'll do our best to help

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Thanks